I don't understand how to solve this question : "It took a boat 5 hours to travel upstream and back. The current was 3 mi/h and the return trip took 1 hour. Find the speed of the boat in still water."
Found 2 solutions by lwsshak3, ankor@dixie-net.com:
Answer by lwsshak3(6490)

I don't understand how to solve this question : "It took a boat 5 hours to travel upstream and back. The current was 3 mi/h and the return trip took 1 hour. Find the speed of the boat in still water.
**
let x=speed of boat in still water
x+3=speed of boat downstream with current
x-3=speed of boat upstream against current
distance=travel time*speed (same for both)
..
5(x-3)=1(x+3)
5x-15=x+3
4x=18
x=18/4=4.5
ans:
speed of boat in still water=4.5 km/hr

Answer by ankor@dixie-net.com(15646)

"It took a boat 5 hours to travel upstream and back.
The current was 3 mi/h and the return trip took 1 hour.
Find the speed of the boat in still water."
:
Let s = boat speed in still water
then
(s-3) = effective speed upstream
and
(s+3) = effective speed down stream
:
It tells us the time up and back was 5 hrs, the return trip (downstream) only took 1 hr, therefore the trip upstream must have taken 4 hrs
:
The distance each way was the same, so write a distance equation, dist = speed * time
4(s-3) = 1(s+3)
4s - 12 = s + 3
4s - s = 3 + 12
3s = 15
s = 15/3
s = 5 mph is the speed of the boat in still water
:
:
Check this by finding the actual distance each way, they should be equal
4(5-3) = 8 mi
1(5+3) = 8 mi
:
:
Did this help you understand this type of problem? Do you have any questions?